OpenStudy (anonymous):

help please!! a 3500 principal earns 3% interest compounded semiannually after 20 years what is the balance of the account? I have never heard of semi annually before? not sure how to solve

OpenStudy (anonymous):

semiannually just means twice a year. so your number of periods is 20 years * 2 = 40 periods

OpenStudy (anonymous):

It would be 7,700 in the end. A = P (1+rt) A = 3,500 (1+.03*40) = 7,700

OpenStudy (anonymous):

thankyou so much